# Almost mathematics of pointed symmetric monoidal model categories by Smith ideal theory

### v4

@@ 4. Bousfield bilocalization of stable model categories
A pointed model category is stable if its homotopy category is a triangulated category. In a stable model category, a homotopy Cartesian square is also homotopy coCartesian and vice versa. Further, in this section, we consider the case that stable monoidal model categories are generated by the monoidal unit objects. Those such categories said to be monogenic, and we explain in Definition 4.1. In this section, we prove that Bousfield bilocalization functors determine homotopically idempotent Smith ideals under the condition that those model categories are stable closed symmetric monoidal and monogenic. Furthermore, we show that homotopically closed objects with respect to those determined Smith ideals by Bousfield bilocalization are homotopically coincide with local objects.
@@ 4.1. Monogenic model categories
Following MacLane [Mac88, Chapter V, Section 7, p.127], we recall the definition of generators of categories: Let $\mathcal{M}$ be a category and $\mathcal{G}$ a class of small objects of $\mathcal{M}$. The set $\mathcal{G}$ is a generator of $\mathcal{M}$ if for any pair of morphisms $h,\,h^{\prime}:X\to Y$ in $\mathcal{M}$, $h\neq h^{\prime}$ implies that there exists $S\in\mathcal{G}$ and a morphism $f:S\to X$ such that $h\circ f\neq h^{\prime}\circ f$. As Hovey’s book [Hov99, p.183, Definition 7.2.1], we define weakly generators of model categories.
Definition 4.1 (c.f. [HPS97] p.7).
@@
By the following proposition, for stable model categories, the above definition coincides with [Hov99, p.183, Definition 7.2.1].
@@
  • proof.
    The if-direction is clear. Let $M$ be an object of $\mathcal{M}$ such that the internal-hom object $\mathrm{Map}_{\mathcal{M}}(S,\,M)$ is contractible for any $S\in\mathcal{G}$. Since $\mathcal{G}$ generates $\mathcal{M}$, for any object $N$ and morphism $f:N\to M$, the condition that $f$ is not null-homotopic implies that there exists an object $S\in\mathcal{G}$ and a morphism $i:S\to N$ and $f\circ i:S\to M$ is not null-homotopic, contradicting the generating condition. $\Box$
@@ 4.2. Symmetric monoidal bilocalization on stable model categories
We consider a Bousfield bilocalization $F:\mathcal{M}\to L(\mathcal{M})$ and let denote $F_{*}$ its fully faithful left adjoint and $U$ its fully faithfully right
${F:\mathcal{M}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces L(\mathcal{M}).}$$\scriptstyle{U}$$\scriptstyle{F_{*}}$
@@
is lax Quillen monoidal. Let $j:I\to V$ be a homotopy image object of $q(V)\circ c_{Q(V)}:F_{*}(F(Q(V))\to Q(V)\to V$. Then all of the canonical morphisms in the diagram
$I\otimes I\rightarrow F_{*}(F(Q(V))\otimes I\rightarrow F_{*}(F(Q(V))\otimes F_{*}(F(Q(V))\rightarrow F_{*}(F(Q(V)))\leftarrow I$
@@
  • proof.
    Let $q(V):Q(V)\to V$ be a cofibrant replacement of $V$. Since the Bousfield localization $F$ is a left and right Quillen functor, $F(Q(V))\to F(V)$ is a trivial fibration, and $F(Q(V))$ is a cofibrant object of $L(\mathcal{M})$. Furthermore, since $F_{*}$ is lax Quillen monoidal, $F(Q(V))$ is equivalent to a monoidal unit of $L(\mathcal{M})$. Therefore, all morphisms in
@@
    are weak equivalences by Ken Brown’s lemma [Hov99, p.6 Lemma 1.1.12]. In particular,
@@
    is a weak equivalence between cofibrant objects. Since $F_{*}$ is a left Quillen functor, by Ken Brown’s lemma [Hov99, p.6 Lemma 1.1.12] again,
@@
    is homotopic to the pushout of $\nabla:F_{*}(F(Q(V)))\otimes F_{*}(F(Q(V))\to F_{*}(F(Q(V))\otimes F(Q(V)))$ along the morphism $F_{*}(F(\mu)):F_{*}(F(Q(V))\otimes F(Q(V)))\to F_{*}(F(Q(V)))$. By the left properness of $\mathcal{M}$ implies that $\mu:F_{*}(F(Q(V)))\otimes F_{*}(F(Q(V)))\to F_{*}(F(Q(V)))$ is a weak equivalence.
By the stability of $\mathcal{M}$, the canonical morphism $(j:I\to V)\to(q(V)\circ c_{Q(V)}:F_{*}(F(Q(V)))\to V)$ is a weak equivalence in the projective model structure. Since $I\to F_{*}(F(Q(V)))$ is a weak equivalence of cofibrant objects, by Ken Brown’s lemma again, both morphisms in $I\otimes I\rightarrow F_{*}(F(Q(V))\otimes I\rightarrow F_{*}(F(Q(V))\otimes F_{*}(F(Q(V))$ are weak equivalences. $\Box$
@@
is lax Quillen monoidal. Then the counit $c:F_{*}\circ F\to\mathrm{Id}_{\mathcal{M}}$ induces determines a homotopically idempotent commutative Smith ideal $j:I\to V$ of a monoidal unit object $V$, where $I$ is a cofibrant replacement of the homotopy image of $q(V)\circ c_{Q(V)}:F_{*}(F(Q(V))\to Q(V)\to V$.
  • proof.
@@
Let $\mathcal{M}$ be a pointed closed symmetric monoidal model with a monoidal unit $V$ and $F:\mathcal{M}\to\mathcal{M}$ a symmetric monoidal Bousfield bilocalization, and $F_{*}:\mathcal{M}\to\mathcal{M}$ denote the homotopically fully faithful left adjoint. Assume that the Quillen adjunction
@@
Let $\mathcal{M}$ be a stable closed symmetric monoidal model category and $F:\mathcal{M}\to L(\mathcal{M})$ a Bousfield bilocalization, and $F_{*}:\mathcal{M}\to\mathcal{M}$ denote the homotopically fully faithful left adjoint. Assume that $\mathcal{M}$ is monogenic and the Quillen adjunction
@@
is lax Quillen monoidal. Let $j:I\to V$ be a homotopically idempotent commutative Smith ideal determined as a result of Proposition 4.5. Let $U$ denote a right Quillen adjoint of $F$. Then all of the morphisms of the diagram
[…]
@@
  • proof.
    Let $j:\tilde{I}\to V$. For any fibrant object $M$ of $\mathcal{M}$, one has a commutative diagram
[…]
@@ 5.1. Operadic Smith ideals
The definitions of colors, symmetric sequences, and colored operads in symmetric monoidal categories can be found in Yau’s book, specifically in [Yau16, Chapter 9 and 11]. Colored operads and maps colored operads are defined in [Yau16, p.176, Definition 11.2.1] and [Yau16, p.180, Definition 11.2.12], respectively. Additionally, the material of algebras over colored operads is explained in [Yau16, p.217, Definition 13.2.1].
Further, in the case that the monoidal structure on $\mathcal{M}$ is closed, algebras over $\mathcal{O}$ are determined by operad maps from $\mathcal{O}$ to endomorphism operads [Yau16, p.226, Definition 13.8.1]. For any $\mathfrak{C}$-colored object $A=\{A_{c}\}_{c\in\mathfrak{C}}$ of $\mathcal{M}$, the endomorphism operad $\mathbf{End}_{\mathcal{M}}(A)$ is defined by
@@
Giving a map $\theta:\mathcal{O}\to\mathbf{End}_{\mathcal{M}}(A)$ of $\mathfrak{C}$-colored operads determines an $\mathcal{O}$-algebra structure on $A$. (See [Yau16, p.231, Proposition 13.9.1]).
@@
Let $\mathcal{M}$ be a symmetric monoidal closed (model) category and $\mathcal{O}$ a colored operad in $\mathcal{M}$. Then we write $\overrightarrow{\mathcal{O}^{\Box}}=L_{1}(\mathcal{O})$ and $\overrightarrow{\mathcal{O}^{\otimes}}=L_{0}(\mathcal{O})$. Since $L_{1}$ (resp. $L_{0}$) is strict monoidal (See [WY24, Definition 3.1.1 (7)]), $\overrightarrow{\mathcal{O}^{\Box}}$ (resp. $\overrightarrow{\mathcal{O}^{\otimes}}$) is a colored operad in $\mathrm{Ar}^{\Box}(\mathcal{M})$ (resp. $\mathrm{Ar}^{\otimes}(\mathcal{M})$). Let $\mathrm{Alg}_{\overrightarrow{\mathcal{O}^{\Box}}}(\mathrm{Ar}^{\Box}(\mathcal{M}))$ (resp. $\mathrm{Alg}_{\overrightarrow{\mathcal{O}^{\otimes}}}(\mathrm{Ar}^{\otimes}(\mathcal{M}))$) denote the category of $\mathcal{O}^{\Box}$-algebra objects of $\mathrm{Ar}^{\Box}(\mathcal{M})$ (resp. $\mathcal{O}^{\otimes}$-algebra objects of $\mathrm{Ar}^{\otimes}(\mathcal{M})$).
Operadic Smith ideals defined as the following:
Definition 5.3 ([WY24] Definition 3.3.1).
@@ 5.2. Non-unital algebras from the (operadic) Smith ideal theory viewpoint
We recall a brief review of non-unital rings, which are rings without assuming the existence of the multiplicative unit $1$. Any unital ring is regarded as a non-unital ring by forgetting the existence of $1$. An augmentation $\varepsilon_{A}:A\to V$ is a ring homomorphism that the unit homomorphism $V\to A$ is a section of $\varepsilon_{A}$, and a $V$-algebra with an augmentation is called an augmented $V$-algebra. Let $\mathrm{Alg}_{V//V}$ denotes the category of augmented $V$-algebras and $\mathrm{Alg}^{\rm nu}_{V}$ the category of non-unital $V$-algebras. It is well-known that there exists a categorical equivalence between $\mathrm{Alg}^{\rm nu}_{V}$ and $\mathrm{Alg}_{V//V}$ as follows: For any $V$-algebra $A$, the direct sum $V\oplus A$ has a canonical unital ring structure defined by
@@
for $m,\,n\in V$ and $a,\,b\in A$. It is easily checked that the functor the adjunction
@@
From the Smith ideal theory viewpoint, the above categorical equivalence is a special case of cokernel-and-kernel correspondence between the arrow categories. We consider the case abelian categories. Let $\mathcal{M}$ be a symmetric closed monoidal abelian category and $\mathrm{Ar}^{\rm im}(\mathcal{M})$ denote the full subcategory spanned those objects $f:R\to S$ such that the unit $f\to\mathrm{ker}(\mathrm{cok}(f))$ is an isomorphism. Then the cokernel functor $\mathrm{Ar}^{\rm im}(\mathcal{M})\to\mathrm{Ar}^{\rm coim}(\mathcal{M})$ is a categorical equivalence.
@@
Let $\mathcal{M}$ be a locally presentable abelian symmetric monoidal category. Then, the arrow category $\mathrm{Ar}(\mathcal{M})$ is also locally presentable and there exist reflective localization functors $L_{\rm im}:\mathrm{Ar}^{\Box}(\mathcal{M})\to\mathrm{Ar}^{\rm im}(\mathcal{M})$ by the unit $\mathrm{id}\to\ker\circ\mathrm{coker}$ and $L_{\rm coim}:\mathrm{Ar}^{\otimes}(\mathcal{M})\to\mathrm{Ar}^{\rm coim}(\mathcal{M})$ by the counit $\mathrm{coker}\circ\ker\to\mathrm{id}$ such that the adjunction
@@
  • proof.
    Since any abelian category is binormal, $f\to\mathrm{im}(f)$ is an isomorphism if and only if $f$ is a monomorphism. By Adámek–Rosický [AR94, p.44, Corollary 1.5.4], the arrow category is locally presentable, and it admits reflective localization. By the definitions of those functors: $\mathrm{im}=\mathrm{ker}\circ\mathrm{cok}$ and $\mathrm{coim}=\mathrm{cok}\circ\mathrm{ker}$, the restriction $\mathrm{cok}:\mathrm{Ar}^{\rm im}(\mathcal{M})\to\mathrm{Ar}^{\rm coim}(\mathcal{M})$ are quasi-inverse functors. $\Box$
@@
Let $\mathcal{M}$ be a locally presentable symmetric monoidal abelian category with a monoidal unit $V$. A commutative non-unital monoid object of $\mathcal{M}$ is a commutative Smith ideal $j:I\to R$ in the localized full subcategory $\mathrm{Ar}^{\rm im}(\mathcal{M})$ satisfying that the cokernel $\mathrm{cok}(j):R\to\mathrm{Coker}(j)$ is isomorphic with an argumentation $\varepsilon_{R}:R\to V$, and $\mathrm{Alg}_{\mathcal{O}}^{\rm nu}(\mathcal{M})$ denote the full subcategory of $\mathrm{Alg}_{\mathcal{O}^{\Box}}(\mathrm{Ar}^{\Box}(\mathcal{M}))$ spanned by non-unital $\mathcal{O}^{\Box}$-algebra objects. In particular, $\mathrm{CAlg}^{\rm nu}(\mathcal{M})$ denotes the full subcategory $\mathrm{CAlg}(\mathrm{Ar}^{\Box}(\mathcal{M}))$ spanned by commutative Smith ideals whose cokernels are isomorphic to $V$. We say that an object of $\mathrm{CAlg}^{\rm nu}(\mathcal{M})$ is a non-unital commutative $V$-algebra object of $\mathcal{M}$.
@@
Let $\mathcal{M}$ be a locally presentable symmetric monoidal abelian category with a monoidal unit $V$. The categorical equivalences

### v5

@@ 4. Bousfield bilocalization of stable model categories
A pointed model category is stable if its homotopy category is a triangulated category. In a stable model category, a homotopy Cartesian square is also homotopy coCartesian and vice versa. Further, in this section, we consider stable monoidal model categories generated by their monoidal unit objects. Such categories are called monogenic, as explained in Definition 4.1. In this section, we prove that Bousfield bilocalization functors determine homotopically idempotent Smith ideals under the condition that the model categories are stable, closed symmetric monoidal, and monogenic. Furthermore, we show that homotopically closed objects with respect to the Smith ideals determined by Bousfield bilocalization homotopically coincide with local objects.
@@ 4.1. Monogenic model categories
Following MacLane [11, Chapter V, Section 7, p.127], we recall the definition of generators of categories: Let $\mathcal{M}$ be a category and $\mathcal{G}$ a class of small objects of $\mathcal{M}$. The set $\mathcal{G}$ is a generator of $\mathcal{M}$ if for any pair of morphisms $h,\,h^{\prime}:X\to Y$ in $\mathcal{M}$, $h\neq h^{\prime}$ implies that there exists $S\in\mathcal{G}$ and a morphism $f:S\to X$ such that $h\circ f\neq h^{\prime}\circ f$. Following Hovey’s book [8, p.183, Definition 7.2.1], we define weak generators of model categories.
Definition 4.1 (c.f. Hovey [8]).
@@
By the following proposition, for stable model categories, the above definition coincides with [8, p.183, Definition 7.2.1].
@@
The if-direction is clear. Let $M$ be an object of $\mathcal{M}$ such that the internal-hom object $\mathrm{Map}_{\mathcal{M}}(S,\,M)$ is contractible for any $S\in\mathcal{G}$. Since $\mathcal{G}$ generates $\mathcal{M}$, for any object $N$ and morphism $f:N\to M$, the condition that $f$ is not null-homotopic implies that there is an object $S\in\mathcal{G}$ and a morphism $i:S\to N$ such that $f\circ i:S\to M$ is not null-homotopic, contradicting the generating condition. $\Box$
@@ 4.2. Symmetric monoidal bilocalization on stable model categories
We consider a Bousfield bilocalization $F:\mathcal{M}\to L(\mathcal{M})$ and let $F_{*}$ denote its fully faithful left adjoint and $U$ its fully faithful right adjoint, respectively.
[…]
@@
is lax Quillen monoidal. Let $j:I\to V$ be a homotopy image object of $q(V)\circ c_{Q(V)}:F_{*}(F(Q(V)))\to Q(V)\to V$. Then all of the canonical morphisms in the diagram
$I\otimes I\rightarrow F_{*}(F(Q(V)))\otimes I\rightarrow F_{*}(F(Q(V)))\otimes F_{*}(F(Q(V)))\rightarrow F_{*}(F(Q(V)))\leftarrow I$
@@
Let $q(V):Q(V)\to V$ be a cofibrant replacement of $V$. Since the Bousfield localization $F$ is both a left and a right Quillen functor, $F(Q(V))\to F(V)$ is a trivial fibration, and $F(Q(V))$ is a cofibrant object of $L(\mathcal{M})$. Furthermore, since $F_{*}$ is lax Quillen monoidal, $F(Q(V))$ is equivalent to a monoidal unit of $L(\mathcal{M})$. Therefore, all morphisms in
@@
are weak equivalences by Ken Brown’s lemma [8, p.6 Lemma 1.1.12]. In particular,
@@
is a weak equivalence between cofibrant objects. Since $F_{*}$ is a left Quillen functor, by Ken Brown’s lemma [8, p.6 Lemma 1.1.12] again,
@@
is homotopic to the pushout of $\nabla:F_{*}(F(Q(V)))\otimes F_{*}(F(Q(V)))\to F_{*}(F(Q(V))\otimes F(Q(V)))$ along the morphism $F_{*}(F(\mu)):F_{*}(F(Q(V))\otimes F(Q(V)))\to F_{*}(F(Q(V)))$. By the left properness of $\mathcal{M}$, $\mu:F_{*}(F(Q(V)))\otimes F_{*}(F(Q(V)))\to F_{*}(F(Q(V)))$ is a weak equivalence.
By the stability of $\mathcal{M}$, the morphism $(j:I\to V)\to(q(V)\circ c_{Q(V)}:F_{*}(F(Q(V)))\to V)$ is a weak equivalence in the projective model structure. Since $I\to F_{*}(F(Q(V)))$ is a weak equivalence of cofibrant objects, by Ken Brown’s lemma again, both morphisms in $I\otimes I\rightarrow F_{*}(F(Q(V)))\otimes I\rightarrow F_{*}(F(Q(V)))\otimes F_{*}(F(Q(V)))$ are weak equivalences. $\Box$
@@
Let $F:\mathcal{M}\to L(\mathcal{M})$ be a Bousfield bilocalization of a stable symmetric monoidal model category, and $F_{*}$ denote its fully faithful left adjoint. Assume that the Quillen adjunction
@@
is lax Quillen monoidal. Then the counit $c:F_{*}\circ F\to\mathrm{Id}_{\mathcal{M}}$ determines a homotopically idempotent commutative Smith ideal $j:I\to V$ of a monoidal unit object $V$, where $I$ is a cofibrant replacement of the homotopy image of $q(V)\circ c_{Q(V)}:F_{*}(F(Q(V)))\to Q(V)\to V$.
@@
Let $\mathcal{M}$ be a pointed closed symmetric monoidal model category with a monoidal unit $V$ and $F:\mathcal{M}\to L(\mathcal{M})$ a symmetric monoidal Bousfield bilocalization, and $F_{*}:L(\mathcal{M})\to\mathcal{M}$ denote the homotopically fully faithful left adjoint. Assume that the Quillen adjunction
@@
Let $\mathcal{M}$ be a stable closed symmetric monoidal model category and $F:\mathcal{M}\to L(\mathcal{M})$ a Bousfield bilocalization, and $F_{*}:L(\mathcal{M})\to\mathcal{M}$ denote the homotopically fully faithful left adjoint. Assume that $\mathcal{M}$ is monogenic and the Quillen adjunction
@@
is lax Quillen monoidal. Let $j:I\to V$ be a homotopically idempotent commutative Smith ideal obtained in Proposition 4.5. Let $U$ denote a right Quillen adjoint of $F$. Then all morphisms in the diagram
$UF(M)\overset{\varphi}{\rightarrow}UF\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,M)\overset{\psi}{\leftarrow}\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,M)$
@@
Let $j:I\to V$. For any fibrant object $M$ of $\mathcal{M}$, one has a commutative diagram
${\mathrm{Hom}_{\mathrm{Ho}(\mathcal{M})}(V,\,UF(M))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\hskip-14.22636pt\varphi_{*}}$${\mathrm{Hom}_{\mathrm{Ho}(\mathcal{M})}(V,\,UF\mathrm{Map}_{\mathcal{M}}(I,\,M))}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces\mathrm{Hom}_{\mathrm{Ho}(\mathcal{M})}(V,\,\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,M))}$$\scriptstyle{\hskip 14.22636pt\psi_{*}}$${\mathrm{Hom}_{\mathrm{Ho}(\mathcal{M})}(F_{*}(F(Q(V))),\,M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Hom}_{\mathrm{Ho}(\mathcal{M})}(F_{*}(F(Q(V)))\otimes I,\,M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\mathrm{Hom}_{\mathrm{Ho}(\mathcal{M})}(\tilde{I},\,M),}$
@@ 5.1. Operadic Smith ideals
The definitions of colors, symmetric sequences, and colored operads in symmetric monoidal categories can be found in standard references on operads; see, for example, White–Yau [18, Section 3].
Further, in the case that the monoidal structure on $\mathcal{M}$ is closed, algebras over $\mathcal{O}$ are determined by operad maps from $\mathcal{O}$ to endomorphism operads. For any $\mathfrak{C}$-colored object $A=\{A_{c}\}_{c\in\mathfrak{C}}$ of $\mathcal{M}$, the endomorphism operad $\mathbf{End}_{\mathcal{M}}(A)$ is defined by
@@
Giving a map $\theta:\mathcal{O}\to\mathbf{End}_{\mathcal{M}}(A)$ of $\mathfrak{C}$-colored operads determines an $\mathcal{O}$-algebra structure on $A$.
@@
Let $\mathcal{M}$ be a symmetric monoidal closed (model) category and $\mathcal{O}$ a colored operad in $\mathcal{M}$. Then we write $\overrightarrow{\mathcal{O}^{\Box}}=L_{1}(\mathcal{O})$ and $\overrightarrow{\mathcal{O}^{\otimes}}=L_{0}(\mathcal{O})$. Since $L_{1}$ (resp. $L_{0}$) is strict monoidal (See [18, Definition 3.1.1 (7)]), $\overrightarrow{\mathcal{O}^{\Box}}$ (resp. $\overrightarrow{\mathcal{O}^{\otimes}}$) is a colored operad in $\mathrm{Ar}^{\Box}(\mathcal{M})$ (resp. $\mathrm{Ar}^{\otimes}(\mathcal{M})$). Let $\mathrm{Alg}_{\overrightarrow{\mathcal{O}^{\Box}}}(\mathrm{Ar}^{\Box}(\mathcal{M}))$ (resp. $\mathrm{Alg}_{\overrightarrow{\mathcal{O}^{\otimes}}}(\mathrm{Ar}^{\otimes}(\mathcal{M}))$) denote the category of $\mathcal{O}^{\Box}$-algebra objects of $\mathrm{Ar}^{\Box}(\mathcal{M})$ (resp. $\mathcal{O}^{\otimes}$-algebra objects of $\mathrm{Ar}^{\otimes}(\mathcal{M})$).
Operadic Smith ideals are defined as follows:
Definition 5.3 ([18] Definition 3.3.1).
@@ 5.2. Non-unital algebras from the (operadic) Smith ideal theory viewpoint
We briefly review non-unital rings, which are rings not assumed to have a multiplicative unit $1$. Any unital ring is regarded as a non-unital ring by forgetting the existence of $1$. An augmentation $\varepsilon_{A}:A\to V$ is a ring homomorphism such that the unit homomorphism $V\to A$ is a section of $\varepsilon_{A}$, and a $V$-algebra with an augmentation is called an augmented $V$-algebra. Let $\mathrm{Alg}_{V//V}$ denote the category of augmented $V$-algebras and $\mathrm{Alg}^{\rm nu}_{V}$ the category of non-unital $V$-algebras. It is well-known that there exists a categorical equivalence between $\mathrm{Alg}^{\rm nu}_{V}$ and $\mathrm{Alg}_{V//V}$ as follows: For any $V$-algebra AA, the direct sum $V\oplus A$ has a canonical unital ring structure defined by
@@
for $m,\,n\in V$ and $a,\,b\in A$. It is easily checked that the adjunction
@@
From the Smith ideal theory viewpoint, the above categorical equivalence is a special case of cokernel-and-kernel correspondence between the arrow categories. We consider the case of abelian categories. Let $\mathcal{M}$ be a symmetric closed monoidal abelian category and $\mathrm{Ar}^{\rm im}(\mathcal{M})$ denote the full subcategory spanned by those objects $f:R\to S$ such that the unit $f\to\mathrm{ker}(\mathrm{cok}(f))$ is an isomorphism. Then the cokernel functor $\mathrm{Ar}^{\rm im}(\mathcal{M})\to\mathrm{Ar}^{\rm coim}(\mathcal{M})$ is a categorical equivalence.
@@
Let $\mathcal{M}$ be a locally presentable abelian symmetric monoidal category. Then the arrow category $\mathrm{Ar}(\mathcal{M})$ is also locally presentable and there exist reflective localization functors $L_{\rm im}:\mathrm{Ar}^{\Box}(\mathcal{M})\to\mathrm{Ar}^{\rm im}(\mathcal{M})$ by the unit $\mathrm{id}\to\ker\circ\mathrm{coker}$ and $L_{\rm coim}:\mathrm{Ar}^{\otimes}(\mathcal{M})\to\mathrm{Ar}^{\rm coim}(\mathcal{M})$ by the counit $\mathrm{coker}\circ\ker\to\mathrm{id}$ such that the adjunction
@@
Since any abelian category is binormal, $f\to\mathrm{im}(f)$ is an isomorphism if and only if $f$ is a monomorphism. By Adámek–Rosický [1, p.44, Corollary 1.5.4], the arrow category is locally presentable, and it admits reflective localization. By the definitions of those functors: $\mathrm{im}=\mathrm{ker}\circ\mathrm{cok}$ and $\mathrm{coim}=\mathrm{cok}\circ\mathrm{ker}$, the restrictions of $\mathrm{cok}$ and $\mathrm{ker}$ to $\mathrm{Ar}^{\rm im}(\mathcal{M})$ and $\mathrm{Ar}^{\rm coim}(\mathcal{M})$ are quasi-inverse functors. $\Box$
@@
Let $\mathcal{M}$ be a locally presentable symmetric monoidal abelian category with a monoidal unit $V$. A commutative non-unital monoid object of $\mathcal{M}$ is a commutative Smith ideal $j:I\to R$ in the localized full subcategory $\mathrm{Ar}^{\rm im}(\mathcal{M})$ satisfying that the cokernel $\mathrm{cok}(j):R\to\mathrm{Coker}(j)$ is isomorphic to an augmentation $\varepsilon_{R}:R\to V$, and $\mathrm{Alg}_{\mathcal{O}}^{\rm nu}(\mathcal{M})$ denotes the full subcategory of $\mathrm{Alg}_{\mathcal{O}^{\Box}}(\mathrm{Ar}^{\Box}(\mathcal{M}))$ spanned by non-unital $\mathcal{O}^{\Box}$-algebra objects. In particular, $\mathrm{CAlg}^{\rm nu}(\mathcal{M})$ denotes the full subcategory $\mathrm{CAlg}(\mathrm{Ar}^{\Box}(\mathcal{M}))$ spanned by commutative Smith ideals whose cokernels are isomorphic to $V$. We say that an object of $\mathrm{CAlg}^{\rm nu}(\mathcal{M})$ is a non-unital commutative $V$-algebra object of $\mathcal{M}$.
@@
Let $\mathcal{M}$ be a locally presentable symmetric monoidal abelian category with a monoidal unit $V$. The categorical equivalence